OpenStudy (caerus):

can anyone help me in this determination of integrating factors..ill give medal

OpenStudy (caerus):

y(2x-9y)dx+2x(x-3y)dy=0

OpenStudy (caerus):

@abb0t

OpenStudy (caerus):

i got IF= \[e ^{x^2}\]

OpenStudy (caerus):

how can i find the general solution for this? \[2x^2e ^{x^2}ydx-xe ^{x^2}y^2dx+ye ^{x^2}dx+xe ^{x^2}dy-ye ^{x^2}dy=0\]

OpenStudy (phi):

you have P(x,y) dx + Q(x,y) dy = 0 so do \[ \int_0^x P(x,y) \ dx + \int_0^y Q(x,y) \ dy \]

OpenStudy (caerus):

ops

OpenStudy (phi):

yes, that is what you do if we use M and N instead of P and Q the first step is to do \[\int M \ dx \]